OpenStudy (anonymous):

How do I solve this? 2y=2-4x

OpenStudy (anonymous):

This is where you come in @sjerman1 ahahah

OpenStudy (anonymous):

this time you are using division

OpenStudy (ryan123345):

divide by 2 on both sides

OpenStudy (anonymous):

so you almost have it in the y=mx+b

OpenStudy (anonymous):

ohh, ok.

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 2 }-\frac{ 4x }{ 2 }\]

OpenStudy (anonymous):

if you wanted it to look like your other problems then you can write this as y=-2x+1

OpenStudy (anonymous):

oh, so that's after you divided though? @sjerman1

OpenStudy (anonymous):

yes mam

OpenStudy (anonymous):

after you divide y=2/2-4x/2 you get y=1-2x

OpenStudy (anonymous):

then i just moved the terms around to get it into point slope form, y=mx+b

OpenStudy (anonymous):

oh, thats easy! I can obviously graph it now! Thanks!!

OpenStudy (anonymous):

no problem!